OpenStudy (anonymous):

the largest number that leaves reminder 7 and 8 when divided by 3248 and 4175 respectively is??

OpenStudy (anonymous):

the answer is 463

OpenStudy (anonymous):

plzzz do it

OpenStudy (anonymous):

i posted this question once and no one gave the steps so i reposted it

OpenStudy (anonymous):

plzz i think chand got it

OpenStudy (anonymous):

Please don't start a new thread about this question. This is already been discussed ad nauseam in that earlier thread and we have explained it you why this question is not correct.

OpenStudy (anonymous):

hey question is not wrong............

OpenStudy (anonymous):

u say it is wrong becozz u cant solve it

OpenStudy (anonymous):

according to the problem, 3248 = xy +7 and 4175 = xz +8 where "x" is the divisor and y,z are the respective quotients. divide both the equations. you get y/z = 7/9

OpenStudy (anonymous):

are you clear until now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@chand: He thinks that you got it, you might like to explain your standpoint.

OpenStudy (anonymous):

yes i knw i suggested this method earlier u se..

OpenStudy (anonymous):

chand plzz go on

OpenStudy (anonymous):

so, let y = 7 and z = 8. this satisfies the equation and gives us x= 463

OpenStudy (anonymous):

Pheww :/ I am out of this thread.

OpenStudy (anonymous):

how on dividing we get y/z=7/9

OpenStudy (anonymous):

xy = 3241 and xz = 4167. divide these

OpenStudy (anonymous):

ok then how y = 7 and z = 8

OpenStudy (anonymous):

basically the question you typed is wrong. you said, " when the number is divided by 3248 and 4175" it should've been " when 3248 and 4175 are divided by " let y = 7k and z=8k. the only value of k satisfying the equation is k=1. so we take y=7 and z=8

OpenStudy (anonymous):

ok thzzz plzzz help http://openstudy.com/study#/updates/4fad2c5be4b059b524f8b3f7

OpenStudy (anonymous):

oh sorry i typed it wrong

OpenStudy (anonymous):

@FoolForMath hope the question now makes some sense.

OpenStudy (anonymous):

It's still the not the largest number.